OpenStudy (dancergirl45):

NEED HELP FAST WILL MEDAL AND FAN!! Given that f(x) = 2x − 5, find the value of x that makes f(x) = 15.

OpenStudy (dancergirl45):

@YoungStudier

OpenStudy (dancergirl45):

Choices are 5 10 20 25

OpenStudy (anonymous):

2 x ___ - 5 = 15

OpenStudy (dancergirl45):

It is either b or d

OpenStudy (anonymous):

Think about it.

OpenStudy (anonymous):

2(5)= ??

OpenStudy (dancergirl45):

yeah I know that but do you go like this f(15)=2x-5 which equals 10 or fx=2*15-5=25

OpenStudy (dancergirl45):

which equation

OpenStudy (anonymous):

I would just solve in your head. You said the choices you think are correct is b and d so for B 2(10)-5=?

OpenStudy (dancergirl45):

15

OpenStudy (anonymous):

Mhmmmm c:

OpenStudy (dancergirl45):

but i already know it is 15

OpenStudy (dancergirl45):

I just dont know which equation to do because they both lead to right answers

OpenStudy (anonymous):

The "x" in f(x) just means what number is filled in for the x in the equation.

OpenStudy (dancergirl45):

ok so the 2nd one and the answer is 25

OpenStudy (anonymous):

it would be B as 2 x 10 = 20 20 - 5 = 15

OpenStudy (anonymous):

Let me check and make sure I am correct as I haven't done this in a very long time.

OpenStudy (dancergirl45):

ok

OpenStudy (dancergirl45):

because you fill x in with 15 so 2*15 is 30 -5=25

OpenStudy (anonymous):

You need the equation on the right of the equal sign to equal 15, So if you put 10 in for x in f(x) then you get 15.

OpenStudy (dancergirl45):

ok thank you

OpenStudy (anonymous):

No problem c: